OpenStudy (inkyvoyd):

2.2: 1 (solving problems)

OpenStudy (inkyvoyd):

http://prntscr.com/9vcsnq

OpenStudy (inkyvoyd):

Solve y'=x^2/y (seperable)

OpenStudy (inkyvoyd):

dy/dx =x^2/y y dy=x^2 dx int y dy = x^2 dx y^2/2=x^3/3+C

OpenStudy (bahrom7893):

you forgot the int on the other side but whatever, doesn't matter lol

OpenStudy (inkyvoyd):

y^2=2/3 x^3+C \(\Huge y=\pm \sqrt{2/3 x^3+C}\) and oh whoops